### v1

@@ 2.7 Least Square Quadratic Twin Support Vector Machine
These analogous models have closed-from solutions. For $k=1$ and $2$, let us introduce the following matrices:
@@ 3 Imbalanced Universum Quadratic Twin Support Vector Machines
Class imbalance in binary classification poses significant challenges, leading to biased models and compromised predictive performance, particularly for minority classes. This imbalance can result in reduced sensitivity, misinterpretation of accuracy, and skewed decision boundaries, with severe consequences in critical domains like fraud detection or medical diagnoses. To address this, we draw inspiration from moosaei2023inverse ; richhariya2020reduced , incorporating Universum points to support minority classes in linearly separable datasets. Extending this to quadratic twin support vector machines, we leverage its flexibility to enhance Universum point effectiveness. Through this approach, we aim to improve classifier robustness and generalization, especially in imbalanced scenarios.
@@
In the following parts, we maintain a general approach where the minority class ($A$) is considered as the positive class. To address the class imbalance, we employ random undersampling of the negative class ($B$) to create a balanced dataset in forming a quadratic surface for the minority class. Precisely, letting $A=\{x_{1},x_{2},\dots,x_{|\mathcal{I}_{1}|}\}$ and $B=\{\bar{x}_{1},\bar{x}_{2},\dots,\bar{x}_{|\mathcal{I}_{2}|}\}$ with $|\mathcal{I}_{1}|\ll|\mathcal{I}_{2}|$, we randomly select a reduced sample $\tilde{B}=\{\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{|\mathcal{I}_{1}|}\}$ from the negative class. We also construct $U=\{u_{1},u_{2},\dots,u_{r}\}$ where $r=|\mathcal{I}_{2}|-|\mathcal{I}_{1}|$ Universum points by averaging technique. Next, we choose a reduced sample of them and construct the reduced Universum sample $\hat{U}:=\{\hat{u}_{1},\hat{u}_{2},\dots,\hat{u}_{g}\}$ with $g=\lceil{|\mathcal{I}_{1}|}/{2}\rceil$. Note that $g\ll r$, and these values are deliberately chosen to ensure that the optimization problems associated with each class remain unbiased when an appropriate number of Universum points is added.
@@
The main model of the paper presented next is designed to handle class imbalance when the positive class is assumed to be the minority. In this Imbalanced Universum Quadratic Twin Support Vector Machine (Im-$\mathfrak{U}$-QTSVM) model, we incorporate only $g$ Universum points in the formulation of the minority class to better refine the generalization boundaries in its favor. To address the bias caused by the majority class, we use $r$ Universum points in its formulation, almost as many as the number of points in this class. Slack variables are also used for penalizing misclassification in the class points and noise in the Universum points. Furthermore, an $\ell_{2}$ regularization term on the Hessian of the quadratic surface has been added to enhance the stability and generalization capability of the model. Consequently, the optimization problem for the Im-$\mathfrak{U}$-QTSVM model, which is based on the hinge loss function, can be formulated as follows:
$\min$ $\frac{1}{2}\sum_{i\in\mathcal{I}_{1}}(\frac{1}{2}{x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+{c_{1}}})^{2}+\frac{1}{2}C_{1}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{1}}_{i}+\frac{1}{2}C_{\hat{u}}\sum_{j=1}^{g}{\psi_{1}}_{j}+\frac{1}{2}\lambda_{1}\sum_{i\leq j}{W_{1}}^{2}_{ij}$
@@
$\min$ $\frac{1}{2}\sum_{i\in\mathcal{I}_{2}}(\frac{1}{2}{\bar{x}_{i}^{T}W_{2}\bar{x}_{i}+b_{2}^{T}\bar{x}_{i}+{c_{2}}})^{2}+\frac{1}{2}C_{2}\sum_{i=1}^{|\mathcal{I}_{2}|}{\xi_{2}}_{i}+\frac{1}{2}C_{u}\sum_{j=1}^{r}{\psi_{2}}_{j}+\frac{1}{2}\lambda_{2}\sum_{i\leq j}{W_{2}}^{2}_{ij}$
@@
Similarly, the least squares version of the model discussed above can manage class imbalance when the positive class is the minority. In the Imbalanced Least Square Universum Quadratic Twin Support Vector Machine (Im-LS-$\mathfrak{U}$-QTSVM) model, we also utilize $g$ Universum points in the minority class formulation to better refine the generalization boundaries in its favor. Similarly, to mitigate the bias caused by the majority class, we employ $r$ Universum points in its formulation, which is nearly equivalent to the number of points in this class. Slack variables are incorporated to penalize misclassification in the class points and noise in the Universum points as well. Additionally, an $\ell_{2}$ regularization term is added to the Hessian of the quadratic surface to improve the stability and generalization capability of the model. Further, we consider the quadratic function as the loss function. Therefore, the optimization problem for the Im-LS-$\mathfrak{U}$-QTSVM model can be expressed as follows:
$\min$ $\frac{1}{2}\sum_{i\in\mathcal{I}_{1}}(\frac{1}{2}{x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+{c_{1}}})^{2}+\frac{1}{2}C_{1}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{1}}^{2}_{i}+\frac{1}{2}C_{\hat{u}}\sum_{j=1}^{g}{\psi_{1}}^{2}_{j}+\frac{1}{2}\lambda_{1}\sum_{i\leq j}{W_{1}}^{2}_{ij}$
@@
$\min$ $\frac{1}{2}\sum_{i\in\mathcal{I}_{2}}(\frac{1}{2}{\bar{x}_{i}^{T}W_{2}\bar{x}_{i}+b_{2}^{T}\bar{x}_{i}+{c_{2}}})^{2}+\frac{1}{2}C_{2}\sum_{j=1}^{|\mathcal{I}_{2}|}{\xi_{2}}^{2}_{i}+\frac{1}{2}C_{u}\sum_{j=1}^{r}{\psi_{2}}^{2}_{j}+\frac{1}{2}\lambda_{2}\sum_{i\leq j}{W_{2}}^{2}_{ij}$
@@
The constraints of the above least square version model are equations, so by substituting $\xi$ and $\psi$ into the objective function, and via Definition 1 and letting
@@
$\qquad+\frac{C_{\hat{u}}}{2}\big{\|}(-1+\varepsilon)e-({S^{T}_{\hat{U}}}z_{1}+c_{1}e)\big{\|}^{2}_{2}+\frac{\lambda_{1}}{2}\|Vz_{1}\|^{2}_{2}.$ (Im-LS-$\mathfrak{U}$-QTSVM-1)
@@
$\left[\begin{matrix}S_{A}S_{A}^{T}+C_{1}{S_{\tilde{B}}}{S^{T}_{\tilde{B}}}+C_{\hat{u}}{S_{\hat{U}}}{S^{T}_{\hat{U}}}+\lambda_{1}V^{T}V&Ae+C_{1}{S_{\tilde{B}}}e+C_{\hat{u}}{S_{\hat{U}}}e\\ e^{T}{S^{T}_{A}}+C_{1}e^{T}{S^{T}_{\tilde{B}}}+C_{\hat{u}}e^{T}{S^{T}_{\hat{U}}}&C_{1}e^{T}e+C_{\hat{u}}e^{T}e+|\mathcal{I}_{1}|\end{matrix}\right]\left[\begin{matrix}z_{1}\\ c_{1}\end{matrix}\right]\vspace{0.9cm}$
@@
Similarly, the problem corresponding to the majority class in the least square loss function becomes:
@@
$+\frac{C_{u}}{2}\big{\|}(1-\varepsilon)e-(S_{U}^{T}z_{2}+c_{2}e)\big{\|}^{2}_{2}+\frac{\lambda_{2}}{2}\|Vz_{2}\|^{2}_{2},$ (Im-LS-$\mathfrak{U}$-QTSVM-2)
which by the same methodology leads to
@@ 4 Theoretical Properties of Proposed Models
Problem (3) has always an optimal solution.
@@
for which the objective function vanishes. If the objective function is zero, then we can deduce that $\xi_{1}=0$, $\psi_{1}=0$ and $W_{1}=0$. Hence the remaining conditions take the form of
@@
This characterizes linear dependence of vectors
@@
or, equivalently, affine dependence of the vectors in the formulation of the theorem. ∎
Naturally, the situation in which $z^{*}=0$ for an optimal solution $(z^{*},c^{*})$ is not desirable. Thus, we present a sufficient condition ensuring that the optimal solution is nonzero. This condition assumes that certain three vectors in space $\mathbb{R}^{\frac{n(n+1)}{2}}$ are linearly independent, which is very likely to be satisfied.
@@
Suppose to the contrary that $z^{*}=0$. From equations (1) we obtain that
@@ 5 Numerical Experiments
In this section, a large scale of numerical experiments are conducted to validate the performance of the proposed Im-LS-$\mathfrak{U}$-QTSVM model for binary classification. We first introduce the experiment settings, including the data sources, the parameter setups, etc. Then all the numerical experiments are conducted to validate the classification accuracy of the proposed model along with some related binary classification models on some public benchmark data sets. Certain statistical tests are also conducted to analyze the computational results.
@@ 5.1 Experiment Settings
In the numerical experiments, multiple benchmark SVM models are implemented for comparison, including the linear SVM, linear Twin SVM and least square twin SVM models. We also tested their corresponding Universum variants. All the tested models and their abbreviations are listed in Table 1. In addition, we list the Python toolboxes (commercial solvers, packages, etc.) in Table 1. All the experiments are conducted on a personal computer with eight Intel (R) Core (TM) i7-13900 @ 3.4 GHz CPUs and 64GB RAM.
@@
Least Square Twin SVM LS-TSVM Numpy $C_{1},C_{2}$
Quadratic Least Square Twin SVM LS-QTSVM Numpy $C_{1},C_{2}$
@@
Least Square Universum Twin SVM LS-$\mathfrak{U}$-TSVM Numpy $C_{1},C_{2},C_{u},\varepsilon$
@@
Least Square Uni. Quadratic Twin SVM LS-$\mathfrak{U}$-QTSVM Numpy $C_{1},C_{2},C_{u},\varepsilon$
Proposed model Im-LS-$\mathfrak{U}$-QTSVM Numpy $C_{1},C_{2},C_{u},\lambda,\varepsilon$
For each experiment, a five-fold cross-validation procedure is applied. Each experiment is repeated ten times for each tested model in order to keep the results statistically meaningful. Means and standard deviations of the accuracy scores are recorded as the measurement of classification effectiveness. Ten percent of data samples from each class is randomly selected to generate Universum data by averaging each pair of samples from the two classes moosaei2023universum .
All the parameters for each model are tuned by applying the grid-search, which is commonly adopted for tuning SVM parameters gao2021novel ; luo2016soft . For each pair of the twin models in TSVM, LS-TSVM, LS-QTSVM, 2.3, LS-$\mathfrak{U}$-TSVM, 2.8, and LS-$\mathfrak{U}$-QTSVM, the penalty parameters $C_{1}$ and $C_{2}$ are selected to be the identical.
We first generate four artificial datasets to show the flexibility of the proposed Im-LS-$\mathfrak{U}$-QTSVM model. The surfaces produced by the proposed model are visualized in Figure 1, in which Figures 1(a) and 1(b) have the similar nonlinear pattern while Figures 1(c) and 1(d) have the other similar one. To show the proposed model on different levels of imbalanced data, we set the imbalanced rates in Figure 1(a) and 1(c) to be three while the imbalanced rates in Figure 1(b) and 1(d) to be ten. The major class of data is in color green while the minor class is in red. For the major class, the number of solid points is the same as the minor class.

### v2

@@ 2.5 Least Squares Quadratic Twin Support Vector Machine
These analogous models have closed-form solutions. For $k=1$ and $2$, let us introduce the following matrices:
@@ 3 Imbalanced Universum Quadratic Twin Support Vector Machines
Class imbalance in binary classification poses significant challenges, leading to biased models and compromised predictive performance, particularly for minority classes. This imbalance can result in reduced sensitivity, misinterpretation of accuracy, and skewed decision boundaries, with severe consequences in critical domains like fraud detection or medical diagnoses. To address this, we draw inspiration from moosaei2023inverse ; richhariya2020reduced , incorporating Universum points to support minority classes in linearly separable datasets. Extending this to quadratic twin support vector machines, we leverage their flexibility to enhance the effectiveness of Universum points. Through this approach, we aim to improve classifier robustness and generalization, especially in imbalanced scenarios.
@@
In the following parts, we maintain a general approach in which the minority class ($A$) is considered the positive class. To address the class imbalance, we employ random undersampling of the negative class ($B$) to create a balanced dataset for forming a quadratic surface for the minority class. Precisely, letting $A=\{x_{1},x_{2},\dots,x_{|\mathcal{I}_{1}|}\}$ and $B=\{\bar{x}_{1},\bar{x}_{2},\dots,\bar{x}_{|\mathcal{I}_{2}|}\}$ with $|\mathcal{I}_{1}|\ll|\mathcal{I}_{2}|$, we randomly select a reduced sample $\tilde{B}=\{\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{|\mathcal{I}_{1}|}\}$ from the negative class. We also construct $U=\{u_{1},u_{2},\dots,u_{r}\}$ where $r=|\mathcal{I}_{2}|-|\mathcal{I}_{1}|$ Universum points using an averaging technique. Next, we choose a reduced sample of them and construct the reduced Universum sample $\hat{U}:=\{\hat{u}_{1},\hat{u}_{2},\dots,\hat{u}_{g}\}$ with $g=\lceil{|\mathcal{I}_{1}|}/{2}\rceil$. Note that $g\ll r$, and these values are deliberately chosen to ensure that the optimization problems associated with each class remain unbiased when an appropriate number of Universum points is added.
@@
The main model of the paper presented next is designed to handle class imbalance when the positive class is assumed to be the minority. In this Imbalanced Universum Quadratic Twin Support Vector Machine (Im-$\mathfrak{U}$-QTSVM) model, we incorporate only $g$ Universum points in the formulation of the minority class to better refine the generalization boundaries in its favor. To address the bias caused by the majority class, we use $r$ Universum points in its formulation, almost as many as the number of points in this class. Slack variables are also used to penalize misclassification in the class points and noise in the Universum points. Furthermore, an $\ell_{2}$ regularization term on the Hessian of the quadratic surface has been added to enhance the stability and generalization capability of the model. Consequently, the optimization problem for the Im-$\mathfrak{U}$-QTSVM model, which is based on the hinge loss function, can be formulated as follows:
$\min$ $\frac{1}{2}\sum_{i\in\mathcal{I}_{1}}(\frac{1}{2}{x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+{c_{1}}})^{2}+\frac{1}{2}C_{1}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{1}}_{i}+\frac{1}{2}C_{\hat{u}}\sum_{j=1}^{g}{\psi_{1}}_{j}+\frac{1}{2}\lambda_{1}\sum_{i\leq j}(W_{1})^{2}_{ij}$
@@
$\min$ $\frac{1}{2}\sum_{i\in\mathcal{I}_{2}}(\frac{1}{2}{\bar{x}_{i}^{T}W_{2}\bar{x}_{i}+b_{2}^{T}\bar{x}_{i}+{c_{2}}})^{2}+\frac{1}{2}C_{2}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{2}}_{i}+\frac{1}{2}C_{u}\sum_{j=1}^{r}{\psi_{2}}_{j}+\frac{1}{2}\lambda_{2}\sum_{i\leq j}(W_{2})^{2}_{ij}$
@@
Similarly, the least squares version of the model discussed above can manage class imbalance when the positive class is the minority. In the Imbalanced Least Squares Universum Quadratic Twin Support Vector Machine (Im-LS-$\mathfrak{U}$-QTSVM) model, we also utilize $g$ Universum points in the minority class formulation to better refine the generalization boundaries in its favor. Similarly, to mitigate the bias caused by the majority class, we employ $r$ Universum points in its formulation, which is nearly equivalent to the number of points in this class. Slack variables are incorporated to penalize misclassification in the class points and noise in the Universum points as well. Additionally, an $\ell_{2}$ regularization term is added to the Hessian of the quadratic surface to improve the stability and generalization capability of the model. Further, we consider the quadratic function as the loss function. Therefore, the optimization problem for the Im-LS-$\mathfrak{U}$-QTSVM model can be expressed as follows:
$\min$ $\frac{1}{2}\sum_{i\in\mathcal{I}_{1}}(\frac{1}{2}{x_{i}^{T}W_{1}x_{i}+b_{1}^{T}x_{i}+{c_{1}}})^{2}+\frac{1}{2}C_{1}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{1}}^{2}_{i}+\frac{1}{2}C_{\hat{u}}\sum_{j=1}^{g}{\psi_{1}}^{2}_{j}+\frac{1}{2}\lambda_{1}\sum_{i\leq j}(W_{1})^{2}_{ij}$
@@
$\min$ $\frac{1}{2}\sum_{i\in\mathcal{I}_{2}}(\frac{1}{2}{\bar{x}_{i}^{T}W_{2}\bar{x}_{i}+b_{2}^{T}\bar{x}_{i}+{c_{2}}})^{2}+\frac{1}{2}C_{2}\sum_{i=1}^{|\mathcal{I}_{1}|}{\xi_{2}}^{2}_{i}+\frac{1}{2}C_{u}\sum_{j=1}^{r}{\psi_{2}}^{2}_{j}+\frac{1}{2}\lambda_{2}\sum_{i\leq j}(W_{2})^{2}_{ij}$
@@
The constraints of the above least squares version model are equations, so by substituting $\xi$ and $\psi$ into the objective function, and via Definition 1 and letting
@@
$\qquad+\frac{C_{\hat{u}}}{2}\big\|(-1+\varepsilon)e-({S^{T}_{\hat{U}}}z_{1}+c_{1}e)\big\|^{2}_{2}+\frac{\lambda_{1}}{2}\|Vz_{1}\|^{2}_{2}.$
@@
$\left[\begin{matrix}S_{A}S_{A}^{T}+C_{1}{S_{\tilde{B}}}{S^{T}_{\tilde{B}}}+C_{\hat{u}}{S_{\hat{U}}}{S^{T}_{\hat{U}}}+\lambda_{1}V^{T}V&Ae+C_{1}{S_{\tilde{B}}}e+C_{\hat{u}}{S_{\hat{U}}}e\\ e^{T}{S^{T}_{A}}+C_{1}e^{T}{S^{T}_{\tilde{B}}}+C_{\hat{u}}e^{T}{S^{T}_{\hat{U}}}&C_{1}e^{T}e+C_{\hat{u}}e^{T}e+|\mathcal{I}_{1}|\end{matrix}\right]\left[\begin{matrix}z_{1}\\ c_{1}\end{matrix}\right]\vskip 25.6073pt$
@@
Similarly, the problem corresponding to the majority class in the least squares loss function becomes:
@@
$+\frac{C_{u}}{2}\big\|(1-\varepsilon)e-(S_{U}^{T}z_{2}+c_{2}e)\big\|^{2}_{2}+\frac{\lambda_{2}}{2}\|Vz_{2}\|^{2}_{2},$
which, by the same methodology, leads to
@@
Remark 1.
For the least squares variant, the closed-form solution guarantees global optimality. For the QP variant, standard convex optimization theory ensures convergence to a global optimum given the convexity of the objective.
Algorithm 1 Imbalanced Universum Quadratic Twin SVM (Im-$\mathfrak{U}$-QTSVM)
Input:
  • Training sets $A$ (minority class), $B$ (majority class)
  • Regularization and penalty parameters $C_{1},C_{2},C_{u},C_{\hat{u}},\lambda_{1},\lambda_{2}$
  • Universum parameter $\varepsilon$
Output: Quadratic surfaces $(W_{1},b_{1},c_{1})$ for the minority class and $(W_{2},b_{2},c_{2})$ for the majority class
  1. 1.
    Randomly undersample the majority class $B$ to obtain a balanced subset $\tilde{B}$ of size $|\mathcal{I}_{1}|$.
  2. 2.
    Construct Universum points:
    • $U=\{u_{1},\dots,u_{r}\}$, where $r=|\mathcal{I}_{2}|-|\mathcal{I}_{1}|$, representing Universum points for the majority class.
    • $\hat{U}=\{\hat{u}_{1},\dots,\hat{u}_{g}\}$, where $g=\lceil|\mathcal{I}_{1}|/2\rceil$, representing Universum points for the minority class.
  3. 3.
    Solve the least-squares optimization problems (3) by solving the corresponding systems of linear equations for each class, and obtain the classifier parameters: (W_1, b_1, c_1)  for the minority class, and  (W_2, b_2, c_2)  for the majority class.
  4. 4.
    Classify a new sample $x\in\mathbb{R}^{n}$ by computing the distances to both quadratic surfaces:
    $\text{Class }k=\arg\min_{i=1,2}\frac{\big|\frac{1}{2}x^{T}W_{i}x+b_{i}^{T}x+c_{i}\big|}{\|W_{i}x+b_{i}\|_{2}^{2}}.$
@@ 4 Theoretical Properties of Proposed Models
The problem (3) always has an optimal solution.
@@
for which the objective function vanishes. If the objective function is zero, then we can deduce that $\xi_{1}=0$, $\psi_{1}=0$ and $W_{1}=0$. Hence, the remaining conditions take the form of
@@
This characterizes the linear dependence of vectors
@@
or, equivalently, the affine dependence of the vectors in the formulation of the theorem. ∎
Naturally, the situation in which $z^{*}=0$ for an optimal solution $(z^{*},c^{*})$ is not desirable. Thus, we present a sufficient condition that ensures the optimal solution is nonzero. This condition assumes that three specific vectors in space $\mathbb{R}^{\frac{n(n+1)}{2}}$ are linearly independent, which is highly likely to be satisfied.
@@
Suppose, on the contrary, $z^{*}=0$. From equations (1) we obtain that
@@
Remark 2(Computational Complexity).
The model (Im-$\mathfrak{U}$-QTSVM) is based on solving two convex quadratic optimization problems. For the class $i\in\{1,2\}$, it uses $\mathcal{O}(n^{2})$ variables $W_{i},b_{i},c_{i}$ and $\mathcal{O}(|\mathcal{I}_{i}|)$ slack variables $\xi_{i},\psi_{i}$. Employing a suitable interior-point method, the resulting complexity is $\mathcal{O}((n^{2}+|\mathcal{I}_{i}|)^{3})$ for class $i$.
For the model (Im-LS-$\mathfrak{U}$-QTSVM), the least squares formulation allows a closed-form solution via two linear systems of equations $\Sigma z=\beta$ and $\tilde{\Sigma}z=\tilde{\beta}$. The constraint matrices require some calculations. The matrix $S_{A}$ has a size of order $n^{2}\times|\mathcal{I}_{1}|$, whence the matrix $S_{A}S_{A}^{T}$ is constructed in time $\mathcal{O}(n^{4}|\mathcal{I}_{1}|)$. Similarly, the construction of matrices $S_{\tilde{B}}S_{\tilde{B}}^{T}$ and $S_{\hat{U}}S_{\hat{U}}^{T}$ needs $\mathcal{O}(n^{4}|\mathcal{I}_{1}|)$ and $\mathcal{O}(n^{4}g)$, respectively. Since the value of $g$ is of order $\mathcal{O}(|\mathcal{I}_{1}|)$, the construction of matrix $\Sigma$ requires a total time of $\mathcal{O}(n^{4}|\mathcal{I}_{1}|)$. The matrix $\Sigma$ has a size of $n^{2}\times n^{2}$, so solving the system of linear equations by the standard methods requires time $\mathcal{O}(n^{6})$. Therefore, the total computational time for the first system of equations is $\mathcal{O}(n^{4}(n^{2}+|\mathcal{I}_{1}|))$. For the second system of equations, we obtain the time complexity $\mathcal{O}(n^{4}(n^{2}+|\mathcal{I}_{2}|))$ analogously.
Hence, the least squares variant is more efficient for large datasets, in particular when $|\mathcal{I}_{i}|\gg n^{2}$.
@@ 5 Numerical Experiments
In this section, a large scale of numerical experiments are conducted to validate the performance of the proposed Im-LS-$\mathfrak{U}$-QTSVM model for binary classification. We first introduce the experimental settings, including the data sources, the parameter setups, etc. Then, all the numerical experiments are conducted to validate the classification accuracy of the proposed model, along with some related binary classification models on public benchmark datasets. Certain statistical tests are also conducted to analyze the computational results.
@@ 5.1 Experiment Settings
In the numerical experiments, multiple benchmark SVM models are implemented for comparison, including the linear SVM, linear Twin SVM, and least squares twin SVM models. We also tested their corresponding Universum variants. All the tested models and their abbreviations are listed in Table 1. In addition, we list the Python toolboxes (commercial solvers, packages, etc.) in Table 1. All experiments are conducted on a MacBook Pro with an Apple M3 Pro chip and 18 GB of unified memory.
@@
Least Squares Twin SVM LS-TSVM Numpy $C_{1},C_{2}$
Quadratic Least Squares Twin SVM LS-QTSVM Numpy $C_{1},C_{2}$
@@
Least Squares Universum Twin SVM LS-$\mathfrak{U}$-TSVM Numpy $C_{1},C_{2},C_{u},\varepsilon$
@@
Least Squares Universum Quadratic Twin SVM LS-$\mathfrak{U}$-QTSVM Numpy $C_{1},C_{2},C_{u},\varepsilon$
Universum Weighted Kernel Extreme Learning Machine URKWELM Numpy $C,C_{u},\gamma$
Intuitionistic Fuzzy Twin Proximal SVM FHTPSVM Numpy $M_{1},M_{2},M_{3},M_{4},$ $\gamma_{1},\gamma_{2},I_{w}$
Imbalanced Least Squares Universum Quadratic Twin SVM Im-LS-$\mathfrak{U}$-QTSVM Numpy $C_{1},C_{2},C_{u},\lambda,\varepsilon$
For each experiment, a five-fold cross-validation procedure is applied. Each experiment is repeated ten times for each tested model to keep the results statistically meaningful. Means and standard deviations of the accuracy scores are recorded as the measurement of classification effectiveness.
Universum Data Generation.
To construct Universum data, we adopt the averaging technique described in moosaei2023universum . The procedure is detailed as follows:
  1. 1.
    Randomly select $10\%$ of samples from each class $A$ and $B$.
  2. 2.
    Pair the selected samples between the two classes.
  3. 3.
    For each pair $(x_{a},x_{b})$, compute a Universum point
    $u=\tfrac{1}{2}(x_{a}+x_{b}).$
  4. 4.
    Collect all such $u$ as the Universum dataset $U$.
All the parameters for each model are tuned by applying the grid-search, which is commonly adopted for tuning SVM parameters gao2021novel ; luo2016soft . For each pair of the twin models in TSVM, LS-TSVM, LS-QTSVM, 2.1, LS-$\mathfrak{U}$-TSVM, 2.6, and LS-$\mathfrak{U}$-QTSVM, the penalty parameters $C_{1}$ and $C_{2}$ are selected to be the identical. In model FHTPSVM, we set $M_{1}=M_{2}$ and $M_{3}=M_{4}$, and $\gamma_{1}=\gamma_{2}=I_{w}=0.5$. In model FHTPSVM, we set kernel parameter $\gamma$ to be the reciprocal of the feature amount. The ranges of the rest of tuning parameters $C$, $C_{u}$, $C_{1}$, $C_{2}$, $M_{1}$, $M_{3}$, $\lambda$ and $\varepsilon$ are searched in grid $\{2^{-8},2^{-7},\dots,2^{8}\}$.
We first generate four artificial datasets to show the flexibility of the proposed Im-LS-$\mathfrak{U}$-QTSVM model. The surfaces produced by the proposed model are visualized in Figure 1, in which Figures 1(a) and 1(b) have the similar nonlinear pattern while Figures 1(c) and 1(d) have the other similar one. To show the proposed model on different levels of imbalanced data, we set the imbalanced rates in Figure 1(a) and 1(c) to be three while the imbalanced rates in Figure 1(b) and 1(d) to be ten. The major class of data is in color green, while the minor class is in red. For the major class, the number of solid points is the same as the minor class.